# Functors on the category of finite sets revisited

### v2

@@ 4.4. A model for the right augmentation to $\mathbbm{k}{\mathbf{FS}}$
(4.4) $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(P^{\mathbf{FA}}_{\mathbf{n}},P^{\mathbf{FA}}_{\bullet})\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\bullet})$
given by applying $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(-,P^{\mathbf{FA}}_{\bullet})$ to (4.3) identifies with the right augmentation (4.2) as a morphism of $\mathbbm{k}\mathfrak{S}_{n}\otimes\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-modules. In particular, there is an isomorphism of $\mathbbm{k}\mathfrak{S}_{n}\otimes\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-modules
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\bullet})\cong\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n}).$
The first part of the argument is isolated in the following Lemma:
Lemma 4.9.
For $n,t\in\mathbb{N}$, there is a commutative diagram
[…]
This induces the morphism of $\mathbbm{k}\mathfrak{S}_{n}\otimes\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-modules
(4.5) $\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n})\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\bullet}).$
@@
It suffices to show that, for $f:\mathbf{t}\rightarrow\mathbf{n}$ a non-surjective map, the image of $[f]$ in $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{t}})$ is zero. Using naturality, by considering the image of $f$, one can reduce to the case $t=n-1$ and $f$ an inclusion $\mathbf{n-1}\hookrightarrow\mathbf{n}$. The result follows in this case, using that $\overline{P}$ is, by definition, the kernel of the map $P^{\mathbf{FA}}_{\mathbf{1}}\rightarrow P^{\mathbf{FA}}_{\mathbf{0}}\cong\mathbbm{k}$. ∎
Proof of Theorem 4.8.
We require to prove that the morphism (4.5) given by Lemma 4.9 is an isomorphism. First we show that it is injective.
Evaluation on $\mathbf{n+1}$ gives the linear map $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{t}})\rightarrow\mathrm{Hom}_{\mathbbm{k}}(\overline{P}^{\otimes n}(\mathbf{n+1}),P^{\mathbf{FA}}_{\mathbf{t}}(\mathbf{n+1})).$ Consider the element $A_{n+1}:=\bigotimes_{i=1}^{n}([i]-[n+1])\in\overline{P}^{\otimes n}(\mathbf{n+1})$; after evaluation, for $\Psi\in\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{t}})$ one obtains $\Psi(A_{n+1})\in P^{\mathbf{FA}}_{\mathbf{t}}(\mathbf{n+1})$, the image of $A_{n+1}$. This yields the $\mathbbm{k}$-linear map
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{t}})$$\rightarrow$ $P^{\mathbf{FA}}_{\mathbf{t}}(\mathbf{n+1})\cong\mathbbm{k}((\mathbf{n+1})^{\times t})\cong(\mathbbm{k}(\mathbf{n+1}))^{\otimes t}$
$\Psi$ $\mapsto$ $\Psi(A_{n+1}).$
Standard arguments imply that this is injective, using that $\{[i]-[n+1]\mid i\in\mathbf{n}\}$ is a basis of $\overline{P}(\mathbf{n+1})$.
In the following, we construct a retract of the map $\mathbbm{k}{\mathbf{FS}}(\mathbf{t},\mathbf{n})\rightarrow\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{n}})$. Using the $t$-fold tensor product of the surjection $\mathbbm{k}(\mathbf{n+1})\twoheadrightarrow\mathbbm{k}\mathbf{n}$ given by sending $[n+1]$ to zero, one has $(\mathbbm{k}(\mathbf{n+1}))^{\otimes t}\twoheadrightarrow(\mathbbm{k}\mathbf{n})^{\otimes t}\cong\mathbbm{k}\mathbf{FA}(\mathbf{t},\mathbf{n})$. (This is not induced by a morphism in $\mathbbm{k}\mathbf{FA}(\mathbf{n+1},\mathbf{n})$.) This gives the composite $\mathbbm{k}$-linear map:
$\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{t}})\rightarrow\mathbbm{k}\mathbf{FA}(\mathbf{t},\mathbf{n}).$
We claim that this maps to $\mathbbm{k}{\mathbf{FS}}(\mathbf{t},\mathbf{n})\subset\mathbbm{k}\mathbf{FA}(\mathbf{t},\mathbf{n})$. Here, under the isomorphism $\mathbbm{k}(\mathbf{n}^{\times t})\cong(\mathbbm{k}\mathbf{n})^{\otimes t}\cong\mathbbm{k}\mathbf{FA}(\mathbf{t},\mathbf{n})$, the subspace $\mathbbm{k}{\mathbf{FS}}(\mathbf{t},\mathbf{n})$ is generated by the tuples in $\mathbf{n}^{\times t}$ in which each element of $\mathbf{n}$ appears at least once.
For each $i\in\mathbf{n}$, one has the self-map $\varphi_{i}$ of $\mathbf{n+1}$ which sends $i$ to $n+1$ and acts as the identity on $\mathbf{n+1}\backslash\{i\}$. Clearly $\varphi_{i}$ sends the element $A_{n+1}\in\overline{P}^{\otimes n}(\mathbf{n+1})$ to zero, hence the same holds for $\Psi(A_{n+1})$ in $P^{\mathbf{FA}}_{\mathbf{t}}(\mathbf{n+1})\cong\mathbbm{k}((\mathbf{n+1})^{\times t})$ for any $\Psi\in\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{t}})$. Suppose that the image of $\Psi(A_{n+1})$ in $\mathbbm{k}(\mathbf{n}^{\times t})\cong\mathbbm{k}\mathbf{FA}(\mathbf{t},\mathbf{n})$ does not lie in $\mathbbm{k}{\mathbf{FS}}(\mathbf{t},\mathbf{n})$; equivalently $\Psi(A_{n+1})=\sum_{\underline{z}\in\mathbf{n+1}^{\times t}}\beta_{\underline{z}}\underline{z}$ where $\beta_{\underline{z}}\neq 0$ for some $\underline{z}\in(\mathbf{n}\backslash\{j\})^{\times t}\subset\mathbf{n+1}^{\times t}$, for some $j\in\mathbf{n}$. From this it follows easily that $\varphi_{j}(\Psi(A_{n+1}))\neq 0$, a contradiction.
Hence this construction yields a $\mathbbm{k}$-linear map
(4.6) $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{t}})\rightarrow\mathbbm{k}{\mathbf{FS}}(\mathbf{t},\mathbf{n}).$
It is straightforward to check that this gives a retract to $\mathbbm{k}{\mathbf{FS}}(\mathbf{t},\mathbf{n})\rightarrow\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{t}})$, which is thus injective.
To conclude, it suffices to show that the map (4.6) is injective. Suppose that $\Psi\in\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{t}})$ is in the kernel of this map. This implies that $\Psi(A_{n+1})\in\mathbbm{k}((\mathbf{n+1})^{\times t})$ can be written as
$\Psi(A_{n+1})=\sum_{\begin{subarray}{c}\underline{x}\in(\mathbf{n+1})^{\times t}\backslash\mathbf{n}^{\times t}\end{subarray}}\lambda_{\underline{x}}\underline{x}.$
The condition on the indexing set means that $n+1$ appears in each tuple $\underline{x}$. Observe that $\Psi$ is non-zero if and only if there exists at least one $\lambda_{\underline{x}}$ that is non-zero.
Consider the inclusion $\iota:\mathbf{n+1}\hookrightarrow\mathbf{n+2}$ given by $\iota(\ell)=\ell$ for $1\leq\ell\leq n$ and $\iota(n+1)=n+2$. The element $\iota(A_{n+1})\in\overline{P}^{\otimes n}(\mathbf{n+2})$ identifies as
$\iota(A_{n+1}):=\bigotimes_{i=1}^{n}([i]-[n+2])$ $=$ $\bigotimes_{i=1}^{n}\big{(}([i]-[n+1])-([n+2]-[n+1])\big{)}$
$=$ $\sum_{Y\subset\mathbf{n}}\bigotimes_{i=1}^{n}(-1)^{|Y|}([\alpha_{Y}(i)]-[n+1]),$
where $\alpha_{Y}(i)$ is $i$ if $i\not\in Y$ and $n+2$ otherwise.
Naturality of $\Psi$ gives the equality $\iota(\Psi(A_{n+1}))=\Psi(\iota(A_{n+1}))$ in $P^{\mathbf{FA}}_{\mathbf{t}}(\mathbf{n+2})$. Clearly one has
$\iota(\Psi(A_{n+1}))=\sum_{\begin{subarray}{c}\underline{x}\in\mathbf{n+1}^{\times t}\backslash\mathbf{n}^{\times t}\end{subarray}}\lambda_{\underline{x}}\iota(\underline{x}),$
where $\iota(\underline{x})$ is given by replacing each occurrence of $n+1$ in $\underline{x}$ by $n+2$. In particular, $n+2$ occurs in the tuple $\iota(\underline{x})$ and $n+1$ does not.
Correspondingly, we have
$\Psi(\iota(A_{n+1}))=\sum_{Y\subset\mathbf{n}}(-1)^{|Y|}\Psi\big{(}\bigotimes_{i=1}^{n}([\alpha_{Y}(i)]-[n+1])),$
where $\alpha_{Y}$ is as above. The map $\alpha_{Y}$ can be extended to a map in $\mathbf{FA}(\mathbf{n+1},\mathbf{n+2})$ by defining $\alpha_{Y}(n+1)=n+1$, so that
$\Psi\big{(}\bigotimes_{i=1}^{n}([\alpha_{Y}(i)]-[n+1])\big{)}=\Psi\big{(}\bigotimes_{i=1}^{n}([\alpha_{Y}(i)]-[\alpha_{Y}(n+1)])\big{)}=\alpha_{Y}(\Psi(A_{n+1})),$
by naturality. The hypothesis on $\Psi(A_{n+1})$ therefore implies that
$\Psi\big{(}\bigotimes_{i=1}^{n}([\alpha_{Y}(i)]-[n+1])\big{)}=\sum_{\underline{z}\in\mathbf{n+2}^{\times t}}\mu^{Y}_{\underline{z}}\underline{z}$
where, if $\mu^{Y}_{\underline{z}}$ is non-zero, then $n+1$ occurs in the tuple $\underline{z}$.
This gives the equality
$\sum_{\begin{subarray}{c}\underline{x}\in\mathbf{n+1}^{\times t}\backslash\mathbf{n}^{\times t}\end{subarray}}\lambda_{\underline{x}}\iota(\underline{x})=\iota(\Psi(A_{n+1}))=\Psi(\iota(A_{n+1}))=\sum_{Y\subset\mathbf{n}}\sum_{\underline{z}}(-1)^{|Y|}\mu^{Y}_{\underline{z}}\underline{z}.$
Given the form of the $\iota(\underline{x})$ and the above condition on the $\underline{z}$ for which $\mu^{Y}_{\underline{z}}\neq 0$, the only way that this equality can hold is if both sides are zero, which implies that $\Psi=0$, as required. ∎
Remark 4.10.
An alternative proof of Theorem 4.8 is as follows. One first calculates $\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(P^{\mathbf{FA}}\otimes\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\mathbf{t}})$ using Corollary 4.3, deducing that there is a short exact sequence of $\mathbbm{k}\mathbf{FA}^{\mathrm{op}}$-modules
(4.7) $0\rightarrow\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n})\rightarrow\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(P^{\mathbf{FA}}\otimes\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\bullet})\rightarrow\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n+1})\rightarrow 0$
One shows that this arises from the short exact sequence
$0\rightarrow\overline{P}^{\otimes n+1}\rightarrow P^{\mathbf{FA}}\otimes\overline{P}^{\otimes n}\rightarrow\overline{P}^{\otimes n}\rightarrow 0$
upon applying the functor $\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(-,P^{\mathbf{FA}}_{\bullet})$.
For this, one considers applying this functor to the composite maps of the form
$P^{\mathbf{FA}}\otimes\overline{P}^{\otimes n+1}\twoheadrightarrow\overline{P}^{\otimes n+1}\hookrightarrow P^{\mathbf{FA}}\otimes\overline{P}^{\otimes n}.$
One analyses the resulting map
$\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(P^{\mathbf{FA}}\otimes\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\bullet})\rightarrow\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(P^{\mathbf{FA}}\otimes\overline{P}^{\otimes n+1},P^{\mathbf{FA}}_{\bullet})$
by exploiting the naturality with respect to $\bullet$ and the respective short exact sequences (4.7). One deduces that the map factorizes across $\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n+1})$ as the composite
$\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(P^{\mathbf{FA}}\otimes\overline{P}^{\otimes n},P^{\mathbf{FA}}_{\bullet})\twoheadrightarrow\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n+1})\hookrightarrow\mathrm{Hom}_{\mathcal{F}(\mathbf{FA})}(P^{\mathbf{FA}}\otimes\overline{P}^{\otimes n+1},P^{\mathbf{FA}}_{\bullet}).$
The key point is that it suffices to analyse the map evaluated on $\bullet=\mathbf{n+1}$, and this can be calculated directly.
@@
$\mathbbm{k}{\mathbf{FS}}(\mathbf{a},\mathbf{n})\rightarrow\bigoplus_{i=1}^{a}\mathbbm{k}{\mathbf{FS}}(\mathbf{a}\backslash\{i\},\mathbf{n}).$
(Here, for a surjection $f:\mathbf{a}\twoheadrightarrow\mathbf{n}$, the $i$th component of the image of $[f]$ is $[f|_{\mathbf{a}\backslash\{i\}}]$, understood to be zero if $f|_{\mathbf{a}\backslash\{i\}}$ is not surjective.)
@@
This follows from Theorem 4.8 by using the the exact sequence $0\rightarrow\overline{P}^{\otimes a}\rightarrow P^{\mathbf{FA}}_{\mathbf{a}}\rightarrow\bigoplus_{i=1}^{a}P^{\mathbf{FA}}_{\mathbf{a}\backslash\{i\}}$ of Proposition 3.5 for $0<a$. ∎
@@
Corollary 4.11 was proved using different methods in [PV23]. The kernel of the total restriction map was then calculated over a general unital commutative ring. A different approach is taken here in Part III, working over a field of characteristic zero.

### v3

@@ 4.3. Morphisms between functors of the form $\mathbf{Q}_{\emptyset}^{X,x}$
induced by the inclusion $\overline{P}^{\otimes n}\subset P^{\mathbf{FA}}_{\mathbf{n}}$ identifies with the right augmentation
$\mathbbm{k}\mathbf{FA}(\bullet,\mathbf{n})\twoheadrightarrow\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n})$
of Proposition 2.11.
@@
The inclusion $\overline{P}^{\otimes n}\subset P^{\mathbf{FA}}_{\mathbf{n}}$ factors across $\mathbf{Q}_{\emptyset}^{\mathbf{n},i}\subset P^{\mathbf{FA}}_{\mathbf{n}}$ for any $i\in\mathbf{n}$. By Corollary 4.24, applying $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(-,P^{\mathbf{FA}}_{\bullet})$ yields surjections
$\mathbbm{k}\mathbf{FA}(\bullet,\mathbf{n})\twoheadrightarrow\mathbf{R}^{\mathbf{n},i}_{\emptyset}(\bullet)\twoheadrightarrow\mathbbm{k}{\mathbf{FS}}(\bullet,\mathbf{n}).$
It remains to check that this composite coincides with the right augmentation. By Yoneda, it suffices to check the image of $\mathrm{Id}_{\mathbf{n}}]\in\mathbbm{k}\mathbf{FA}(\mathbf{n},\mathbf{n})$; this is straightforward. ∎
These results also allow us to recover information on morphisms between tensor products of $\overline{P}$:
@@
[…]
@@
By Proposition 3.7, there is an exact sequence
$0\rightarrow\overline{P}^{\otimes a}\rightarrow P^{\mathbf{FA}}_{\mathbf{a}}\rightarrow\bigoplus_{i=1}^{a}P^{\mathbf{FA}}_{\mathbf{a}\backslash\{i\}}.$
The result follows by applying $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\overline{P}^{\otimes n},-)$ and using Theorem 4.23 to identify the terms and the maps. ∎
@@
Corollary 4.26 was proved using different methods in [7].
4.5. Finiteness properties
The above results imply finiteness properties for certain $\mathbbm{k}\mathbf{FA}$-modules. In preparation, we note the following basic property:
Lemma 4.28.
For a $\mathbbm{k}\mathbf{FA}$-module $F$ and positive integers $m\leq n$, if $F(\mathbf{n})=0$ then $F(\mathbf{m})=0$.
Proof.
The standard inclusion $\mathbf{m}\subset\mathbf{n}$ clearly admits a retract in $\mathbf{FA}(\mathbf{n},\mathbf{m})$, since $\mathbf{m}\neq\emptyset$ by hypothesis. Hence the $\mathbbm{k}$-module $F(\mathbf{m})$ is a direct summand of $F(\mathbf{n})$. ∎
Remark 4.29.
The example of $\mathbbm{k}_{\mathbf{0}}$ shows that the hypothesis upon $m$ is necessary.
We have the following general consequence of Corollary 4.8.
Proposition 4.30.
For $0<t\in\mathbb{N}$ and subfunctors $F\subset G\subset\overline{P}^{\otimes t}$, if $F(\mathbf{t+1})\subset G(\mathbf{t+1})$ is an equality then $F=G$.
Proof.
We require to prove that the quotient $G/F$ is zero. We have $\overline{P}^{\otimes t}(\mathbf{0})=0$, so that the hypothesis implies that $G/F(\mathbf{n})=0$ for each $n\leq t+1$, by Lemma 4.28.
Suppose that $G/F\neq 0$. By the above, we have $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbbm{k},G/F)=0$, thus the set of projective generators given in Corollary 4.8 implies that there exists a pointed finite set $(X,x)$ and a non-zero morphism $\mathbf{Q}^{X,x}_{\emptyset}\rightarrow G/F$. Since $\mathbf{Q}^{X,x}_{\emptyset}$ is projective, this lifts to a non-zero morphism $\mathbf{Q}^{X,x}_{\emptyset}\rightarrow G$; composing with $G\subset\overline{P}^{\otimes t}$ gives a non-trivial morphism $\mathbf{Q}^{X,x}_{\emptyset}\rightarrow\overline{P}^{\otimes t}$. Since $\overline{P}^{\otimes t}$ is a subfunctor of $\mathbf{Q}^{\mathbf{t},t}_{\emptyset}$ (using that $t>0$), by Corollary 4.16, this yields a non-trivial morphism $\mathbf{Q}^{X,x}_{\emptyset}\rightarrow\mathbf{Q}^{\mathbf{t},t}_{\emptyset}$. The vanishing result of Theorem 4.19 therefore implies that $|X|\leq t+1$.
Since $\mathbf{Q}_{\emptyset}^{X,x}$ is a direct summand of $P^{\mathbf{FA}}_{X}$, Yoneda’s lemma gives that $\mathrm{Hom}_{\mathbbm{k}\mathbf{FA}}(\mathbf{Q}_{\emptyset}^{X,x},G/F)$ is a direct summand of $G/F(X)$. However, the latter was shown to be zero, which establishes a contradiction. Thus $G/F=0$, as required. ∎
Recall that an object $A$ in an abelian category is simple (or irreducible) if its only subobjects are 0 and $A$. A similar argument to the above then shows the following:
Proposition 4.31.
Suppose that $F\neq 0$ is a $\mathbbm{k}\mathbf{FA}$-module such that $F(\mathbf{0})=0$. Then there exists $t\in\mathbb{N}$ and a non-trivial morphism $\overline{P}^{\otimes t}\rightarrow F$. In particular, if $F$ is a simple functor, then $F$ is a quotient of $\overline{P}^{\otimes t}$.
If $F(\mathbf{0})\neq 0$, there is a surjection $F\twoheadrightarrow F(\mathbf{0})$ in $\mathcal{F}(\mathbf{FA})$. Hence $F$ is a simple $\mathbbm{k}\mathbf{FA}$-module if and only if $F=F(\mathbf{0})$ and the latter is a simple $\mathbbm{k}$-module.
Proof.
For the first statement, as in the proof of Proposition 4.30, since $F(\mathbf{0})=0$, there exists a finite pointed set $(X,x)$ and a non-zero morphism $\mathbf{Q}^{X,x}_{\emptyset}\rightarrow F$. Using the short exact sequence of Corollary 4.16, either this restricts to a non-zero morphism $\overline{P}^{\otimes X}\rightarrow F$ or it factors across a non-zero morphism $\overline{P}^{\otimes X\backslash\{x\}}\rightarrow F$. The conclusion in the case that $F$ is simple is immediate.
The second statement is straightforward. ∎
Recall that a composition series of an object $A$ in an abelian category is a finite length filtration
$0=A_{0}\subset A_{1}\subset\ldots\subset A_{\ell}=A$
with the property that each $A_{i+1}/A_{i}$ is a simple object. By the Jordan-Hölder theorem, the associated graded $\bigoplus_{i=0}^{\ell-1}A_{i+1}/A_{i}$ is independent (up to isomorphism) of the choice of composition series. The composition factors of $A$ are the simple functors $A_{i+1}/A_{i}$, which can occur with multiplicity greater than one.
Corollary 4.32.
Suppose that $\mathbbm{k}$ is a field and that $t$ is a positive integer. Then $\overline{P}^{\otimes t}$ admits a (finite length) composition series and, for each composition factor $S$ of $\overline{P}^{\otimes t}$:
  1. (1)
    $S(\mathbf{t+1})\neq 0$;
  2. (2)
    $S$ is a quotient of $\overline{P}^{\otimes n}$ for some $n\leq t+1$.
Proof.
The $\mathbbm{k}$-vector space $\overline{P}^{\otimes t}(\mathbf{t+1})$ has finite dimension, so that this follows directly from the two previous Propositions and their proofs. ∎
Using Corollary 4.16, this implies:
Corollary 4.33.
For $\mathbbm{k}$ a field, the indecomposable projectives of $\mathcal{F}(\mathbf{FA})$ admit (finite length) composition series.
Proof.
By Corollary 4.8, an indecomposable projective of $\mathcal{F}(\mathbf{FA})$ is either isomorphic to $P^{\mathbf{FA}}_{\mathbf{0}}=\mathbbm{k}$ or is a direct summand of some $\mathbf{Q}^{\mathbf{n},n}_{\emptyset}$, for $0<n\in\mathbb{N}$. The functor $\mathbbm{k}$ has composition series exhibited by (3.1), using Remark 3.1. In the second case, the result follows from Corollary 4.32 in conjunction with the short exact sequence of Corollary 4.16. ∎